Hello,

I need help with this problem:

An airplane can fly a rate of 525 mi/h in still air. It can fly 2800 miles with the wind in the same time that it can fly 2450 miles against the wind. Find the rate if the wind.

Thank you!

Answer by stanbon(57347)

An airplane can fly a rate of 525 mi/h in still air. It can fly 2800 miles with the wind in the same time that it can fly 2450 miles against the wind. Find the rate of the wind.
--------------
Let the rate of the wind be "w".
---------------
With the wind DATA:
Distance=2800 miles ; Rate = (525+w) mph ; Time = d/r = 2800/(525+w) hrs.
-----------------------------
Against the wind DATA:
Distance=2450 miles : Rate = (525-w) mph ; Time = d/r = 2450/(525-w) hrs.
------------------------------
EQUATION:
time = time
2800/(525+w) = 2450/(525-w)
56/(525+w) = 49/(525-w)
56*525-56w = 49*525+49w
105w = 7*525
w = 35 mph (the speed of the wind is 35 mph)
================
cheers,
Stan H.
---------------------------------